Suppose a person standing at the top of a building 112 ft high throws a ball vertically upward
with an initial velocity of 96 ft/sec.
(a) Find the velocity of the ball at time
1
Expert's answer
2020-11-30T12:19:02-0500

Since the time was not specified (incomplete), and there was no equation given.

Assume the time you meant was 2s.


Using the equation;

v = u + at

At the top of the building,

u = 96ft/sec

a = -g = 32.2ft/s²

t = 2s


v = 96 + (-32)2

v = 96 - 64

V = 32m/s


Therefore the velocity at t = 2s is 32m/s.